# Mind Meets Robots: A Review of EEG-Based Brain-Robot Interaction Systems

### v1

@@ 2.1. Brain Robot Interaction and EEG Signals
Bi et al. 2013 (Bi et al., 2013)
Krishnan et al. 2016 (Krishnan et al., 2016)
Mao et al. 2017 (Mao et al., 2017)
Aljalal et al. 2020 (Aljalal et al., 2020)
Huang et al. 2021 (Huang and Wang, 2021)
@@ 2.2. Current EEG-based BRI Systems
Several reviews have mapped the terrain of EEG-based BRI and associated fields. In 2010, Zhang et al. (Zhang et al., 2010) provided a summary of BCI’s evolution in industrial robotics and evaluated new commercial BCI products. Subsequently, Si-Mohammed et al. (Si-Mohammed et al., 2017) conducted a literature review merging BCI with augmented reality (AR), covering applications in medicine, robotics, home automation, and brain activity visualization. Hwang et al. (Hwang et al., 2013) provided a detailed account of EEG-based BCI studies from 2007-2011, and Cao et al. (Cao, 2020) later updated this landscape in 2020, with a focus on integrating EEG-based BCI and artificial intelligence (AI). In robotics, Bi et al. (Bi et al., 2013) comprehensively reviewed EEG-driven control for mobile robots, addressing systems, techniques, and evaluations, while Krishhan et al. (Krishnan et al., 2016) focused on EEG control for assistive robots aiding the disabled and elderly. Aljalal et al. (Aljalal et al., 2020) surveyed EEG signal processing for robot control, highlighting challenges in noninvasive BCI systems. They highlighted encountered challenges of brain-controlled systems at the time. Huang et al. (Huang and Wang, 2021) delved into EEG signal processing methods, spotlighting neural network and deep learning (DL) techniques for signal classification. In 2017, Mao (Mao et al., 2017) reviewed the research output generated from the past years with a precise emphasis on EEG-based BRI interactive systems, which aligns with the thematic intent of our paper the most so far. They identified the key techniques and BCI paradigms, but without focusing on the interaction between brains and robots. In fact, none of the previously mentioned papers investigated the brain-robot interaction mode, a critical subject in the HRI community. Moreover, our observation indicates that research in this area has notably increased since 2018 (Fig.  3.A), yet reviews synthesizing these recent advances are still missing. While some have then explored EEG-based robot-assisted rehabilitation (Berger et al., 2019), or EEG-based BCI interaction with virtual reality (VR) and AR (Nwagu et al., 2023), there’s an urgent need for a current review that consolidates the latest EEG-based BRI techniques, highlighting the interaction between human brains and robotic systems. Table  1 showcases the extensive advancements our paper offers over prior similar reviews. We’ve integrated in-depth insights on key techniques, robot classification, interaction modes, applications, challenges, and future directions.
@@ 4.1.1. Signal Acquisition
EEG Paradigm
Sensory and motor-related, SMR/ERD
1 (1.1%)
(Nann et al., 2020)
Sensory and motor-related, MRCP
1 (1.1%)
(Sugiyama et al., 2023)
Sensory and motor-related, MI
21 (24.1%)
@@
(Chenga et al., 2023; Tang et al., 2022; Roy and Bhaumik, 2022; Liu and Jebelli, 2021; Martínez-Cagigal et al., 2020; Penaloza et al., 2018; Hernandez-Carmona and Penaloza, 2019; Li et al., 2019; Wei et al., 2021; Araujo et al., 2021; Francis et al., 2021; Xu et al., 2018; Boonarchatong and Ketcham, 2023; Ak et al., 2022; Zhao et al., 2020b; Guo et al., 2020; Gordleeva et al., 2020; Kuffuor and Samanta, 2018; Jiang et al., 2018; Wang et al., 2018b; Ai et al., 2018)
Vision-related, SSVEP
17 (19.5%)
@@
(Farmaki et al., 2022; Chhabra et al., 2020; Fang et al., 2022; Lyu et al., 2022; Si-Mohammed et al., 2018; Chu et al., 2018; Chen et al., 2021; Abougarair et al., 2021; Zhang et al., 2021; Du et al., 2021; Aznan et al., 2019; Chiuzbaian et al., 2019; Shao et al., 2020; Karunasena et al., 2021; Yuan and Li, 2018; Chen et al., 2020; Quiles et al., 2022)
Cognition-related, P300
8 (9.2%)
@@
(Belkacem and Lakas, 2021; Cao and Liu, 2018; Braun et al., 2019; Li et al., 2023; Magee and Givigi, 2021; Bahman and Shamsollahi, 2019; Rahul and Sharma, 2019; Ali et al., 2021)
Cognition-related, ERP
2 (2.2%)
(Aldini et al., 2022; Kompatsiari et al., 2018)
Cognition-related, ErrP
1 (1.1%)
(Iwane et al., 2019)
Cognition-related, task-based
33 (37.9%)
@@
(Baka et al., 2019; Prinsen et al., 2022; Wu et al., 2022; Richter et al., 2023; Memar and Esfahani, 2018; Dissanayake et al., 2022; Chang and Sun, 2021; Aldini et al., 2019; Toichoa Eyam et al., 2021; Staffa and Rossi, 2022; Wang and Sugaya, 2021; Memar and Esfahani, 2019; Lu et al., 2022; Alimardani et al., 2022, 2020; Long et al., 2019; Yoon et al., 2021; Ogino and Mitsukura, 2018; Pawuś and Paszkiel, 2022; Cervantes et al., 2023; Kilmarx et al., 2018; Wang et al., 2018c; Li et al., 2022a; Jo et al., 2022; Lu et al., 2020; Roshdy et al., 2021; Sanguantrakul et al., 2020; Yu et al., 2019; Ghosh and Orlando, 2019; Mondini et al., 2020; Rashid et al., 2020; Korovesis et al., 2019; Ehrlich and Cheng, 2019)
Cognition-related, IAPS
1 (1.1%)
(Kim et al., 2021)
Hybrid
1 (1.1%)
(Kar et al., 2021)
Acquisition Device
ADS1299
wired/wireless
2 (2.2%)
(Li et al., 2022a; Sanguantrakul et al., 2020)
BIOPAC MP150
wired
1 (1.1%)
(Ghosh and Orlando, 2019)
BioRadio system
wireless
1 (1.1%)
(Karunasena et al., 2021)
Biosemi ActiveTwo
wired
3 (3.4%)
(Tang et al., 2022; Long et al., 2019; Lyu et al., 2022)
BrainProducts MOVE
wireless
1 (1.1%)
(Aldini et al., 2022)
BrainProducts ActiChamp
wired
1 (1.1%)
(Ehrlich and Cheng, 2019)
BrainProducts ActiCap
wired/wireless
4 (4.6%)
@@
(Martínez-Cagigal et al., 2020; Kompatsiari et al., 2018; Mondini et al., 2020; Araujo et al., 2021)
BrainProducts LiveAmp
wireless
1 (1.1%)
(Nann et al., 2020)
BrainProducts others
wired
2 (2.2%)
(Kim et al., 2021; Aldini et al., 2019)
Cognionics HD-72
wireless
1 (1.1%)
(Qian et al., 2018)
Cognionics Quick-20
wireless
1 (1.1%)
(Aznan et al., 2019)
EGI system
wired
1 (1.1%)
(Alimardani et al., 2022)
Emotiv EPOC
wireless
19 (21.8%)
@@
(Roshdy et al., 2021; Pawuś and Paszkiel, 2022; Chang and Sun, 2021; Kilmarx et al., 2018; Yu et al., 2019; Bahman and Shamsollahi, 2019; Chu et al., 2018; Chiuzbaian et al., 2019; Chhabra et al., 2020; Staffa and Rossi, 2022; Wang et al., 2018b; Ai et al., 2018; Kuffuor and Samanta, 2018; Lu et al., 2020; Shao et al., 2020; Braun et al., 2019; Li et al., 2023; Ali et al., 2021; Toichoa Eyam et al., 2021)
g.HIAMP
wired
1 (1.1%)
(Richter et al., 2023)
g.MOBIlab+
wireless
1 (1.1%)
(Dissanayake et al., 2022)
g.Tec Nautilus
wired
1 (1.1%)
(Hernandez-Carmona and Penaloza, 2019)
g.USBamp
wired/N.S.
3 (3.4%)
(Yoon et al., 2021; Alimardani et al., 2020; Si-Mohammed et al., 2018)
Neuracle
wired/N.S.
3 (3.4%)
(Chen et al., 2020; Abougarair et al., 2021; Chen et al., 2021)
Neuroelectrics
wireless
1 (1.1%)
(Quiles et al., 2022)
NeuroScan SynAmps
wired
2 (2.2%)
(Fang et al., 2022; Jo et al., 2022)
NeuroScan NuAmps
wireless
4 (4.6%)
(Li et al., 2019; Baka et al., 2019; Guo et al., 2020; Yuan and Li, 2018)
NeuroScan Others
wired/wireless/N.S.
4 (4.6%)
(Magee and Givigi, 2021; Du et al., 2021; Cao and Liu, 2018; Wang et al., 2018c)
NeuroSky Mindwave
wireless/N.S.
6 (6.9%)
@@
(Rahul and Sharma, 2019; Rashid et al., 2020; Wang and Sugaya, 2021; Ogino and Mitsukura, 2018; Ak et al., 2022; Boonarchatong and Ketcham, 2023)
NeuroSky TGAM
wireless
1 (1.1%)
(Lu et al., 2022)
Nexus 32
wireless
1 (1.1%)
(Cervantes et al., 2023)
Nihon Kohden
N.S.
1 (1.1%)
(Kar et al., 2021)
NVX 52
wired
1 (1.1%)
(Gordleeva et al., 2020)
OpenBCI Cyton
wireless
1 (1.1%)
(Farmaki et al., 2022)
OpenBCI Ganglion
N.S.
1 (1.1%)
(Korovesis et al., 2019)
Unicorn Hybrid Black
wireless
2 (2.2%)
(Belkacem and Lakas, 2021; Prinsen et al., 2022)
Sensor Locations
One of F,T,P,O Involved
14 (16.1%)
@@
(Guo et al., 2020; Kar et al., 2021; Belkacem and Lakas, 2021; Farmaki et al., 2022; Chiuzbaian et al., 2019; Memar and Esfahani, 2018; Ghosh and Orlando, 2019; Wang and Sugaya, 2021; Kuffuor and Samanta, 2018; Korovesis et al., 2019; Chenga et al., 2023; Li et al., 2023; Sanguantrakul et al., 2020; Martínez-Cagigal et al., 2020)
Two of F,T,P,O Involved
21 (24.1%)
@@
(Abougarair et al., 2021; Araujo et al., 2021; Karunasena et al., 2021; Zhang et al., 2021; Yuan and Li, 2018; Iwane et al., 2019; Cao and Liu, 2018; Chen et al., 2021; Bahman and Shamsollahi, 2019; Jiang et al., 2018; Rashid et al., 2020; Ogino and Mitsukura, 2018; Dissanayake et al., 2022; Yu et al., 2019; Du et al., 2021; Chen et al., 2020; Si-Mohammed et al., 2018; Kompatsiari et al., 2018; Shao et al., 2020; Lu et al., 2022; Rahul and Sharma, 2019)
Three of F,T,P,O Involved
9 (10.3%)
@@
(Wei et al., 2021; Prinsen et al., 2022; Nann et al., 2020; Fang et al., 2022; Chu et al., 2018; Aznan et al., 2019; Quiles et al., 2022; Magee and Givigi, 2021; Penaloza et al., 2018)
All of F,T,P,O Involved
13 (14.9%)
@@
(Baka et al., 2019; Cervantes et al., 2023; Ak et al., 2022; Ehrlich and Cheng, 2019; Zhao et al., 2020b; Chang and Sun, 2021; Tang et al., 2022; Wang et al., 2018b; Ai et al., 2018; Long et al., 2019; Staffa and Rossi, 2022; Chhabra et al., 2020; Alimardani et al., 2022)
Earlobe (A/M) Involved
21 (24.1%)
@@
(Cervantes et al., 2023; Ak et al., 2022; Zhao et al., 2020b; Wang et al., 2018c, b; Ai et al., 2018; Ogino and Mitsukura, 2018; Long et al., 2019; Staffa and Rossi, 2022; Dissanayake et al., 2022; Chu et al., 2018; Yu et al., 2019; Aznan et al., 2019; Du et al., 2021; Chhabra et al., 2020; Wang and Sugaya, 2021; Kuffuor and Samanta, 2018; Korovesis et al., 2019; Kompatsiari et al., 2018; Rahul and Sharma, 2019; Martínez-Cagigal et al., 2020)
Midline Sagittal Plane (Z) Involved
22 (25.3%)
@@
(Cervantes et al., 2023; Araujo et al., 2021; Yoon et al., 2021; Nann et al., 2020; Karunasena et al., 2021; Zhang et al., 2021; Jiang et al., 2018; Wang et al., 2018b; Rashid et al., 2020; Ai et al., 2018; Li et al., 2019; Staffa and Rossi, 2022; Dissanayake et al., 2022; Aznan et al., 2019; Chen et al., 2020; Korovesis et al., 2019; Kompatsiari et al., 2018; Quiles et al., 2022; Shao et al., 2020; Rahul and Sharma, 2019; Martínez-Cagigal et al., 2020; Penaloza et al., 2018)
Number of Electrodes
No. $\leq$ 10
37 (42.5%)
@@
(Wang and Sugaya, 2021; Xu et al., 2018; Rashid et al., 2020; Ogino and Mitsukura, 2018; Chiuzbaian et al., 2019; Rahul and Sharma, 2019; Wu et al., 2022; Roy and Bhaumik, 2022; Karunasena et al., 2021; Guo et al., 2020; Zhang et al., 2021; Li et al., 2019; Lu et al., 2022; Yuan and Li, 2018; Ghosh and Orlando, 2019; Yu et al., 2019; Korovesis et al., 2019; Sanguantrakul et al., 2020; Ali et al., 2021; Bahman and Shamsollahi, 2019; Iwane et al., 2019; Si-Mohammed et al., 2018; Magee and Givigi, 2021; Gordleeva et al., 2020; Kar et al., 2021; Prinsen et al., 2022; Nann et al., 2020; Fang et al., 2022; Belkacem and Lakas, 2021; Dissanayake et al., 2022; Du et al., 2021; Quiles et al., 2022; Hernandez-Carmona and Penaloza, 2019; Abougarair et al., 2021; Chen et al., 2021, 2020; Kuffuor and Samanta, 2018)
10 $<$ No. $\leq$ 20
23 (26.4%)
@@
(Alimardani et al., 2022; Lu et al., 2020; Roshdy et al., 2021; Braun et al., 2019; Toichoa Eyam et al., 2021; Alimardani et al., 2020; Kilmarx et al., 2018; Yoon et al., 2021; Chang and Sun, 2021; Cao and Liu, 2018; Staffa and Rossi, 2022; Chu et al., 2018; Chhabra et al., 2020; Li et al., 2023; Cervantes et al., 2023; Wei et al., 2021; Araujo et al., 2021; Wang et al., 2018b; Ai et al., 2018; Jiang et al., 2018; Sugiyama et al., 2023; Memar and Esfahani, 2019; Aznan et al., 2019)
20 $<$ No. $\leq$ 32
14 (16.1%)
@@
(Pawuś and Paszkiel, 2022; Ak et al., 2022; Penaloza et al., 2018; Wang et al., 2018c; Kim et al., 2021; Aldini et al., 2019; Jo et al., 2022; Lyu et al., 2022; Aldini et al., 2022; Liu and Jebelli, 2021; Qian et al., 2018; Ehrlich and Cheng, 2019; Long et al., 2019; Shao et al., 2020)
32 $<$ No. $\leq$ 64
5 (5.7%)
(Baka et al., 2019; Martínez-Cagigal et al., 2020; Richter et al., 2023; Tang et al., 2022; Kompatsiari et al., 2018)
No. $>$ 64
3 (3.4%)
(Mondini et al., 2020; Farmaki et al., 2022; Francis et al., 2021)

### v3

@@ 2.1. Brain Robot Interaction and EEG Signals
Bi et al. 2013 (Bi et al., 2013) - - - -
Krishnan et al. 2016 (Krishnan et al., 2016) - - - -
Mao et al. 2017 (Mao et al., 2017) - - -
Aljalal et al. 2020 (Aljalal et al., 2020) - -
Huang et al. 2021 (Huang and Wang, 2021) - - -
@@ 2.2. Current EEG-based BRI Systems
Several reviews have mapped the terrain of EEG-based BRI and associated fields. In 2010, Zhang et al. (Zhang et al., 2010) provided a summary of BCI’s evolution in industrial robotics and evaluated new commercial BCI products. Subsequently, Si-Mohammed et al. (Si-Mohammed et al., 2017) conducted a literature review merging BCI with augmented reality (AR), covering applications in medicine, robotics, home automation, and brain activity visualization. Hwang et al. (Hwang et al., 2013) provided a detailed account of EEG-based BCI studies from 2007-2011, and Cao et al. (Cao, 2020) later updated this landscape in 2020, with a focus on integrating EEG-based BCI and artificial intelligence (AI). In robotics, Bi et al. (Bi et al., 2013) comprehensively reviewed EEG-driven control for mobile robots, addressing systems, techniques, and evaluations, while Krishhan et al. (Krishnan et al., 2016) focused on EEG control for assistive robots aiding the disabled and elderly. Aljalal et al. (Aljalal et al., 2020) surveyed EEG signal processing for robot control, highlighting challenges in noninvasive BCI systems. They highlighted encountered challenges of brain-controlled systems at the time. Huang et al. (Huang and Wang, 2021) delved into EEG signal processing methods, spotlighting neural network and deep learning (DL) techniques for signal classification. In 2017, Mao (Mao et al., 2017) reviewed the research output generated from the past years with a precise emphasis on EEG-based BRI interactive systems, which aligns with the thematic intent of our paper the most so far. They identified the key techniques and BCI paradigms, but without focusing on the interaction between brains and robots. In fact, none of the previously mentioned papers investigated the brain-robot interaction mode, a critical subject in the HRI community. Moreover, our observation indicates that research in this area has notably increased since 2018 (Fig.  3.A), yet reviews synthesizing these recent advances are still missing. While some have then explored EEG-based robot-assisted rehabilitation (Berger et al., 2019), or EEG-based BCI interaction with virtual reality (VR) and AR (Nwagu et al., 2023), there’s an urgent need for a current review that consolidates the latest EEG-based BRI techniques, highlighting the interaction between human brains and robotic systems. Table  LABEL:compa showcases the extensive advancements our paper offers over prior similar reviews. We’ve integrated in-depth insights on key techniques, robot classification, interaction modes, applications, challenges, and future directions.
@@ 4.1.1. Signal Acquisition
EEG Paradigm Sensory and motor-related, SMR/ERD 1 (1.1%) (Nann et al., 2020)
Sensory and motor-related, MRCP 1 (1.1%) (Sugiyama et al., 2023)
@@
Sensory and motor-related, MI 21 (24.1%) (Chenga et al., 2023; Tang et al., 2022; Roy and Bhaumik, 2022; Liu and Jebelli, 2021; Martínez-Cagigal et al., 2020; Penaloza et al., 2018; Hernandez-Carmona and Penaloza, 2019; Li et al., 2019; Wei et al., 2021; Araujo et al., 2021; Francis et al., 2021; Xu et al., 2018; Boonarchatong and Ketcham, 2023; Ak et al., 2022; Zhao et al., 2020b; Guo et al., 2020; Gordleeva et al., 2020; Kuffuor and Samanta, 2018; Jiang et al., 2018; Wang et al., 2018b; Ai et al., 2018)
@@
Vision-related, SSVEP 17 (19.5%) (Farmaki et al., 2022; Chhabra et al., 2020; Fang et al., 2022; Lyu et al., 2022; Si-Mohammed et al., 2018; Chu et al., 2018; Chen et al., 2021; Abougarair et al., 2021; Zhang et al., 2021; Du et al., 2021; Aznan et al., 2019; Chiuzbaian et al., 2019; Shao et al., 2020; Karunasena et al., 2021; Yuan and Li, 2018; Chen et al., 2020; Quiles et al., 2022)
@@
Cognition-related, P300 8 (9.2%) (Belkacem and Lakas, 2021; Cao and Liu, 2018; Braun et al., 2019; Li et al., 2023; Magee and Givigi, 2021; Bahman and Shamsollahi, 2019; Rahul and Sharma, 2019; Ali et al., 2021)
Cognition-related, ERP 2 (2.2%) (Aldini et al., 2022; Kompatsiari et al., 2018)
Cognition-related, ErrP 1 (1.1%) (Iwane et al., 2019)
@@
Cognition-related, task-based 33 (37.9%) (Baka et al., 2019; Prinsen et al., 2022; Wu et al., 2022; Richter et al., 2023; Memar and Esfahani, 2018; Dissanayake et al., 2022; Chang and Sun, 2021; Aldini et al., 2019; Toichoa Eyam et al., 2021; Staffa and Rossi, 2022; Wang and Sugaya, 2021; Memar and Esfahani, 2019; Lu et al., 2022; Alimardani et al., 2022, 2020; Long et al., 2019; Yoon et al., 2021; Ogino and Mitsukura, 2018; Pawuś and Paszkiel, 2022; Cervantes et al., 2023; Kilmarx et al., 2018; Wang et al., 2018c; Li et al., 2022a; Jo et al., 2022; Lu et al., 2020; Roshdy et al., 2021; Sanguantrakul et al., 2020; Yu et al., 2019; Ghosh and Orlando, 2019; Mondini et al., 2020; Rashid et al., 2020; Korovesis et al., 2019; Ehrlich and Cheng, 2019)
Cognition-related, IAPS 1 (1.1%) (Kim et al., 2021)
Hybrid 1 (1.1%) (Kar et al., 2021)
Acquisition Device ADS1299 wired/wireless 2 (2.2%) (Li et al., 2022a; Sanguantrakul et al., 2020)
BIOPAC MP150 wired 1 (1.1%) (Ghosh and Orlando, 2019)
BioRadio system wireless 1 (1.1%) (Karunasena et al., 2021)
Biosemi ActiveTwo wired 3 (3.4%) (Tang et al., 2022; Long et al., 2019; Lyu et al., 2022)
BrainProducts MOVE wireless 1 (1.1%) (Aldini et al., 2022)
BrainProducts ActiChamp wired 1 (1.1%) (Ehrlich and Cheng, 2019)
@@
BrainProducts ActiCap wired/wireless 4 (4.6%) (Martínez-Cagigal et al., 2020; Kompatsiari et al., 2018; Mondini et al., 2020; Araujo et al., 2021)
BrainProducts LiveAmp wireless 1 (1.1%) (Nann et al., 2020)
BrainProducts others wired 2 (2.2%) (Kim et al., 2021; Aldini et al., 2019)
Cognionics HD-72 wireless 1 (1.1%) (Qian et al., 2018)
Cognionics Quick-20 wireless 1 (1.1%) (Aznan et al., 2019)
EGI system wired 1 (1.1%) (Alimardani et al., 2022)
@@
Emotiv EPOC wireless 19 (21.8%) (Roshdy et al., 2021; Pawuś and Paszkiel, 2022; Chang and Sun, 2021; Kilmarx et al., 2018; Yu et al., 2019; Bahman and Shamsollahi, 2019; Chu et al., 2018; Chiuzbaian et al., 2019; Chhabra et al., 2020; Staffa and Rossi, 2022; Wang et al., 2018b; Ai et al., 2018; Kuffuor and Samanta, 2018; Lu et al., 2020; Shao et al., 2020; Braun et al., 2019; Li et al., 2023; Ali et al., 2021; Toichoa Eyam et al., 2021)
g.HIAMP wired 1 (1.1%) (Richter et al., 2023)
g.MOBIlab+ wireless 1 (1.1%) (Dissanayake et al., 2022)
g.Tec Nautilus wired 1 (1.1%) (Hernandez-Carmona and Penaloza, 2019)
g.USBamp wired/N.S. 3 (3.4%) (Yoon et al., 2021; Alimardani et al., 2020; Si-Mohammed et al., 2018)
Neuracle wired/N.S. 3 (3.4%) (Chen et al., 2020; Abougarair et al., 2021; Chen et al., 2021)
Neuroelectrics wireless 1 (1.1%) (Quiles et al., 2022)
NeuroScan SynAmps wired 2 (2.2%) (Fang et al., 2022; Jo et al., 2022)
NeuroScan NuAmps wireless 4 (4.6%) (Li et al., 2019; Baka et al., 2019; Guo et al., 2020; Yuan and Li, 2018)
NeuroScan Others wired/wireless/N.S. 4 (4.6%) (Magee and Givigi, 2021; Du et al., 2021; Cao and Liu, 2018; Wang et al., 2018c)
@@
NeuroSky Mindwave wireless/N.S. 6 (6.9%) (Rahul and Sharma, 2019; Rashid et al., 2020; Wang and Sugaya, 2021; Ogino and Mitsukura, 2018; Ak et al., 2022; Boonarchatong and Ketcham, 2023)
NeuroSky TGAM wireless 1 (1.1%) (Lu et al., 2022)
Nexus 32 wireless 1 (1.1%) (Cervantes et al., 2023)
Nihon Kohden N.S. 1 (1.1%) (Kar et al., 2021)
NVX 52 wired 1 (1.1%) (Gordleeva et al., 2020)
OpenBCI Cyton wireless 1 (1.1%) (Farmaki et al., 2022)
OpenBCI Ganglion N.S. 1 (1.1%) (Korovesis et al., 2019)
Unicorn Hybrid Black wireless 2 (2.2%) (Belkacem and Lakas, 2021; Prinsen et al., 2022)
@@
Sensor Locations One of F,T,P,O Involved 14 (16.1%) (Guo et al., 2020; Kar et al., 2021; Belkacem and Lakas, 2021; Farmaki et al., 2022; Chiuzbaian et al., 2019; Memar and Esfahani, 2018; Ghosh and Orlando, 2019; Wang and Sugaya, 2021; Kuffuor and Samanta, 2018; Korovesis et al., 2019; Chenga et al., 2023; Li et al., 2023; Sanguantrakul et al., 2020; Martínez-Cagigal et al., 2020)
@@
Two of F,T,P,O Involved 21 (24.1%) (Abougarair et al., 2021; Araujo et al., 2021; Karunasena et al., 2021; Zhang et al., 2021; Yuan and Li, 2018; Iwane et al., 2019; Cao and Liu, 2018; Chen et al., 2021; Bahman and Shamsollahi, 2019; Jiang et al., 2018; Rashid et al., 2020; Ogino and Mitsukura, 2018; Dissanayake et al., 2022; Yu et al., 2019; Du et al., 2021; Chen et al., 2020; Si-Mohammed et al., 2018; Kompatsiari et al., 2018; Shao et al., 2020; Lu et al., 2022; Rahul and Sharma, 2019)
@@
Three of F,T,P,O Involved 9 (10.3%) (Wei et al., 2021; Prinsen et al., 2022; Nann et al., 2020; Fang et al., 2022; Chu et al., 2018; Aznan et al., 2019; Quiles et al., 2022; Magee and Givigi, 2021; Penaloza et al., 2018)
@@
All of F,T,P,O Involved 13 (14.9%) (Baka et al., 2019; Cervantes et al., 2023; Ak et al., 2022; Ehrlich and Cheng, 2019; Zhao et al., 2020b; Chang and Sun, 2021; Tang et al., 2022; Wang et al., 2018b; Ai et al., 2018; Long et al., 2019; Staffa and Rossi, 2022; Chhabra et al., 2020; Alimardani et al., 2022)
@@
Earlobe (A/M) Involved 21 (24.1%) (Cervantes et al., 2023; Ak et al., 2022; Zhao et al., 2020b; Wang et al., 2018c, b; Ai et al., 2018; Ogino and Mitsukura, 2018; Long et al., 2019; Staffa and Rossi, 2022; Dissanayake et al., 2022; Chu et al., 2018; Yu et al., 2019; Aznan et al., 2019; Du et al., 2021; Chhabra et al., 2020; Wang and Sugaya, 2021; Kuffuor and Samanta, 2018; Korovesis et al., 2019; Kompatsiari et al., 2018; Rahul and Sharma, 2019; Martínez-Cagigal et al., 2020)
@@
Midline Sagittal Plane (Z) Involved 22 (25.3%) (Cervantes et al., 2023; Araujo et al., 2021; Yoon et al., 2021; Nann et al., 2020; Karunasena et al., 2021; Zhang et al., 2021; Jiang et al., 2018; Wang et al., 2018b; Rashid et al., 2020; Ai et al., 2018; Li et al., 2019; Staffa and Rossi, 2022; Dissanayake et al., 2022; Aznan et al., 2019; Chen et al., 2020; Korovesis et al., 2019; Kompatsiari et al., 2018; Quiles et al., 2022; Shao et al., 2020; Rahul and Sharma, 2019; Martínez-Cagigal et al., 2020; Penaloza et al., 2018)
@@
Number of Electrodes No. $\leq$ 10 37 (42.5%) (Wang and Sugaya, 2021; Xu et al., 2018; Rashid et al., 2020; Ogino and Mitsukura, 2018; Chiuzbaian et al., 2019; Rahul and Sharma, 2019; Wu et al., 2022; Roy and Bhaumik, 2022; Karunasena et al., 2021; Guo et al., 2020; Zhang et al., 2021; Li et al., 2019; Lu et al., 2022; Yuan and Li, 2018; Ghosh and Orlando, 2019; Yu et al., 2019; Korovesis et al., 2019; Sanguantrakul et al., 2020; Ali et al., 2021; Bahman and Shamsollahi, 2019; Iwane et al., 2019; Si-Mohammed et al., 2018; Magee and Givigi, 2021; Gordleeva et al., 2020; Kar et al., 2021; Prinsen et al., 2022; Nann et al., 2020; Fang et al., 2022; Belkacem and Lakas, 2021; Dissanayake et al., 2022; Du et al., 2021; Quiles et al., 2022; Hernandez-Carmona and Penaloza, 2019; Abougarair et al., 2021; Chen et al., 2021, 2020; Kuffuor and Samanta, 2018)
@@
10 $<$ No. $\leq$ 20 23 (26.4%) (Alimardani et al., 2022; Lu et al., 2020; Roshdy et al., 2021; Braun et al., 2019; Toichoa Eyam et al., 2021; Alimardani et al., 2020; Kilmarx et al., 2018; Yoon et al., 2021; Chang and Sun, 2021; Cao and Liu, 2018; Staffa and Rossi, 2022; Chu et al., 2018; Chhabra et al., 2020; Li et al., 2023; Cervantes et al., 2023; Wei et al., 2021; Araujo et al., 2021; Wang et al., 2018b; Ai et al., 2018; Jiang et al., 2018; Sugiyama et al., 2023; Memar and Esfahani, 2019; Aznan et al., 2019)
@@
20 $<$ No. $\leq$ 32 14 (16.1%) (Pawuś and Paszkiel, 2022; Ak et al., 2022; Penaloza et al., 2018; Wang et al., 2018c; Kim et al., 2021; Aldini et al., 2019; Jo et al., 2022; Lyu et al., 2022; Aldini et al., 2022; Liu and Jebelli, 2021; Qian et al., 2018; Ehrlich and Cheng, 2019; Long et al., 2019; Shao et al., 2020)
32 $<$ No. $\leq$ 64 5 (5.7%) (Baka et al., 2019; Martínez-Cagigal et al., 2020; Richter et al., 2023; Tang et al., 2022; Kompatsiari et al., 2018)
No. $>$ 64 3 (3.4%) (Mondini et al., 2020; Farmaki et al., 2022; Francis et al., 2021)